page 7 8. The following is Ameline Corporations Stockholders Equity for the years, 2018 and 2019: Stockholders Equity 2018
Answer #1
2018 2019 Changes
Common stock 2,00,000 2,00,000
Retained earnings 11,50,000 13,98,120 1,398,120 - 1,150,000 = 248,120
Total stockholder equity 13,50,000 15,98,120
Dividend 55,000
a.
Addition to Retained earnings = Net income - Dividend
Thus, Net income = Addition to retained earnings + Dividend
=248,120 + 55,000 = 303,120
So Net income in 2019 is $303,120
b.
Debt ratio is 30% means the remaining value which is equity is 70% of total assets
So we have equity value i.e. 1,350,000 in 2018 and we can find out the total assets
Debt ratio = Total debt / total assets
Equity ratio = Total Equity / total assets
Thus, Total assets = Total Equity / Equity ratio
1,350,000 / 70% = 1,928,571.43
So Total assets in 2018 is 1,928,571.43